# Compact quantum algorithms that can potentially maintain quantum advantage for solving time-dependent differential equations

### v1

@@ Abstract
Many claims of computational advantages have been made for quantum computing over classical, but they have not been demonstrated for practical problems. Here, we present algorithms for solving time-dependent PDEs governing fluid flow problems. We build on an idea based on linear combination of unitaries to simulate non-unitary, non-Hermitian quantum systems, and generate hybrid quantum-classical algorithms that efficiently perform iterative matrix-vector multiplication and matrix inversion operations. These algorithms lead to low-depth quantum circuits that protect quantum advantage, with the best-case asymptotic complexities that are near-optimal. We demonstrate the performance of the algorithms by conducting: (a) ideal state-vector simulations using an in-house, high performance, quantum simulator called QFlowS; (b) experiments on a real quantum device (IBM Cairo); and (c) noisy simulations using Qiskit Aer. We also provide device specifications such as error-rates (noise) and state sampling (measurement) to accurately perform convergent flow simulations on noisy devices.
@@ I Introduction
Simulating nonlinear phenomena is quite arduous. In the case of hydrodynamic turbulence, for instance, the combination of a wide range of scales and the need for fine computational resolution [yeung2020advancing] makes it quite challenging [iyer2021area, buaria2022scaling]. Similar challenges are encountered in other problems such as glassy and molecular dynamics, protein folding and chemical reactions. Simulating such phenomena requires a paradigm shift in computing—and a strong candidate for that is Quantum Computing (QC). QC is successful in surpassing its classical counterparts in some cases [alexeev2021quantum, awschalom2022roadmap], but it is yet to demonstrate its might in solving practical problems. See, e.g., [bharadwaj2020quantum, bharadwajintroduction] for a discussion of problems involving fluid dynamics. A general bottleneck is that most classical systems are nonlinear, whereas quantum algorithms that consist of quantum gates and circuits obey laws of quantum mechanics, which are linear and unitary. For the numerical tools used in this work, reconciling this mismatch makes linearization of the problem inevitable [lin2022koopman, giannakis2022embedding, liu2021efficient], leading to high dimensional linear problems. It is thus necessary to create efficient quantum algorithms to solve high dimensional linear system of equations. Even without linearization, solving nonlinear PDEs would still demand, at the least, an efficient way to iteratively operate general non-unitary and non-hermitian matrices. The current work focuses precisely on developing these tools.
To solve flow problems (and PDEs in general) using QC, different quantum algorithms have already been proposed—e.g., Quantum Linear System Algorithms (QLSA) [harrow2009quantum, childs2017quantum, subacsi2019quantum, liu2021efficient, childs2021high, fang2023time, bharadwaj2023hybrid] and Variational Quantum Algorithms (VQA) [leong2022variational, lubasch2020variational, ingelmann2023two]. These algorithms solve the governing PDEs from a continuum scale approach. There have also been meso-scale approaches based on Lattice Boltzmann methods [todorova2020quantum, budinski2021quantum, itani2024quantum]. A somewhat newer approach called Schrödingerization [jin2022quantum, meng2023quantum, jin2024quantum] uses analog qubits (continuous variables) to map classical PDEs into equivalent quantum systems, and integrates the corresponding Schrödinger equation. Although the present work uses the continuum approach, the same tools can be applied also to the discrete systems and others involving iterative matrix-vector multiplications and inversions. The examples include data analysis, machine learning, image processing, optimization, and so forth.
@@
In this work we propose a set of hybrid, quantum linear systems algorithms based on Linear Combination of Unitaries (LCU) for solving the fluid equations. The current approach can provide provable guarantees on gate and time complexities. We present six Time Marching Compact Quantum Circuit (TMCQC) algorithms to solve the time-dependent PDEs by explicit and implicit time marching schemes. We make use of a certain concept of LCU, which has been used earlier in a different context of non-hermitian open quantum systems [schlimgen2021quantum] and quantum chemistry applications. We transform this tool into a hybrid quantum-classical algorithm that use a time marching approach. These algorithms require only two, or at most four, controlled unitaries for the LCU decomposition leading to low-depth quantum circuits. We show that, of the six TMCQCs proposed, the gate complexity (of basic two-qubit gates) is at best $\mathcal{O}\Big{(}s\epsilon,\log(N_{g}\tau),\textrm{polylog}(\epsilon/\varepsilon_{U},1/\varepsilon_{N}),\log^{-3}((\kappa-1)^{-1})\Big{)}$. Here, $N_{g}$ is the grid size, $\tau$ is the number of time steps, $s,\kappa$ are the sparsity and condition numbers of the matrix operators, and $\epsilon,\varepsilon_{U},\varepsilon_{N}$ determine the accuracy of different approximations of the algorithm, as discussed in later sections.
The above complexity scaling is near-optimal in all parameters except $s$ (which, however, still contributes as a small constant prefactor owing to the sparse, tri-diagonal matrices used here). At the worst, the complexity could be exponential in both $\tau$ and $s$. As one would expect, an LCU tool based on the time marching approach would generally tend to quickly diminish the success of the algorithm, thus blowing up the sampling sizes required to recover the solution accurately. However, using tools such as the Richardson extrapolation, we also show that the query complexity of the algorithm can also be kept to a minimum, such that it contributes only a constant pre-factor to the overall complexity. The qubit complexity scales as $\mathcal{O}(\log(N_{g}\tau))$ at best and $\mathcal{O}(\log(N_{g}\tau(\log(\varepsilon^{-1}_{N}))/\log((\kappa-1)^{-1})))$ at the worst, which is optimal in both cases. Note that the complexities mentioned here are without any quantum amplitude amplification. The algorithms also avoid the need for expensive phase estimation methods, trotterization schemes, bit-arithmetic and classical optimization subroutines, making the quantum circuits compact. Furthermore, we also propose a few end-to-end strategies, using which the overall circuit simulations can conserve any available quantum advantage, apart from also making them amenable on current and near-term hardware.
We use these algorithms to simulate a linear advection-diffusion problem and use its analytical and classical solutions as reference for estimating the accuracy of the solutions obtained. The simple and well-understood nature of the chosen problem makes it an ideal candidate for assessing the performance of the quantum algorithm. However, the tools described here can be extended to a more general class of flow problems. The performance of the current algorithms is studied via (a) statevector simulations using QFlowS—an in-house, high-performance, quantum simulator [bharadwaj2023hybrid], (b) experiments with a real quantum device (IBM Cairo), and (c) noisy simulations with IBM Qiskit Aer platform to study noise effects by separating error contributions due to the algorithm itself from finite sampling of the final quantum states. The results of these three exercises indicate that our algorithms can produce accurate results that capture the flow physics both qualitatively and quantitatively. We use insights from these results to prescribe parameters for circuit designs to perform accurate, forward time simulations of flow problems. We also estimate the resource and specification requirements of near-term devices that would be necessary to perform convergent flow simulations. We further highlight that, although we solve a simplified flow problem, these algorithms are quite general and can be used to solve nonlinear problems as well, by appropriately linearizing the problem into any one of the TMCQC formats, thus broadening their appeal. These results suggest that the proposed algorithms can be simulated on near-term machines at full scale, while retaining a potential quantum advantage.
@@
We remark in passing that even though variational algorithms could be quite efficient in practice, it is hard to prove results rigorously owing to their strong dependence on the underlying optimization algorithm [ingelmann2023two].
@@ II Governing equations
where the velocity varies only along y (wall-normal direction). We set $C=U$ (the constant advection velocity in the x-direction), and we put $D$, representing the diffusion coefficient, to unity. The initial condition $u(L/2,0)$ is chosen as the delta function $\delta(x)$ and the system is subject to a periodic boundary condition $u(0,t)=u(L,t)$. Such a setting also admits an analytical solution [ingelmann2023two], making it an ideal test-case to evaluate the accuracy of the quantum solutions. The algorithmic framework developed here is also agnostic to the linearity or otherwise of the PDEs. For instance, if we were to begin by considering the nonlinear case ($C=\mathbf{u}(x,t)$), a preceding linearization such as a Homotopy Analysis and Carlemann or Koopman method [liu2021efficient, giannakis2022embedding, lin2022koopman, bharadwaj2023quantum] can be applied to first obtain an approximate, higher dimensional linear system of equations. Such a system can then be solved using the proposed methods.
@@ IV.6 Query complexity and success probability
    does not diverge. If the time marching operator is norm-preserving (or the flow is statistically steady), then $||~{}|\Psi\rangle_{\tau}~{}||=||~{}|\Psi\rangle_{0}~{}||$ trivially. The alternative scenario of the norm-increasing flow is better, where $||~{}|\Psi\rangle_{\tau}~{}||\geq||~{}|\Psi\rangle_{0}~{}||$, corresponding to flows that experience constantly increasing net influx of energy. Both cases are in practice possible if we consider a constant forcing term in our governing PDEs. However, the more restrictive case is the viscous, dissipative flow whose norm decays with time; in this case the steady state of the flow corresponds to vanishingly small velocity fields due to constant dissipation of energy from the system. However, the advection-diffusion problem considered in this work is not as severe as can be seen from the analytical solution [ingelmann2023two]. The steady state corresponds to a constant, uniform velocity field with a small, yet non-zero norm (in this work, this ratio is = 1/2). For flows whose steady state norms are identically zero, the query complexity grows as $\mathcal{O}(||~{}|\Psi\rangle_{\tau}~{}||^{-1})$. In such cases, there could be two alternatives (which should be studied in more detail): (1) The problem of the growing pre-factor $\eta$ can be addressed by choosing a smaller $\epsilon$ such that $\eta^{G_{\textrm{L}}}\sim\mathcal{O}(||~{}|\Psi\rangle_{\tau}~{}||^{-1})$. (2) When this choice is not possible, one is limited to solving the problem for only short time horizons, for which the norm decays not too steeply.

### v2

# Compact quantum algorithms for time-dependent differential equations

@@ Abstract
Many claims of computational advantages have been made for quantum computing over classical, but they have not been demonstrated for practical problems. Here, we present algorithms for solving time-dependent PDEs, with particular reference to fluid equations. We build on an idea based on linear combination of unitaries to simulate non-unitary, non-Hermitian quantum systems, and generate hybrid quantum-classical algorithms that efficiently perform iterative matrix-vector multiplication and matrix inversion operations. Though these algorithms are end-to-end, they require relatively low-depth quantum circuits and protect quantum advantage, with the best-case asymptotic complexities, which we show are near-optimal. We demonstrate the performance of the algorithms by conducting: (a) fully gate level, state-vector simulations using an in-house, high performance, quantum simulator called QFlowS; (b) experiments on a real quantum device; and (c) noisy simulations using Qiskit Aer. We also provide device specifications such as error-rates (noise) and state sampling (measurement) to accurately perform convergent flow simulations on noisy devices. The results offer evidence that the proposed algorithm is amenable for use on near-term quantum devices.
@@ I Introduction
Simulating nonlinear phenomena is quite arduous. In the case of hydrodynamic turbulence, for instance, the combination of a wide range of scales and the need for fine computational resolution [yeung2020advancing] makes it quite challenging [iyer2021area, buaria2022scaling]. Similar challenges are encountered in other problems such as glassy and molecular dynamics, protein folding and chemical reactions. Simulating such phenomena requires a paradigm shift in computing—and a strong candidate for it is Quantum Computing (QC). QC may surpass classical digital computing in some cases [alexeev2021quantum, awschalom2022roadmap], but it is yet to demonstrate its power in solving practical problems. See, e.g., Refs. [bharadwaj2020quantum, bharadwajintroduction, bharadwaj2024simulating] for a discussion of this point, especially on fluid dynamics. A general point is that most classical systems are nonlinear, whereas quantum algorithms—consisting of quantum gates and quantum circuits—obey laws of quantum mechanics, which are linear and unitary. Reconciling this mismatch makes linearization of the problem inevitable [lin2022koopman, giannakis2022embedding, liu2021efficient], leading to high dimensional linear problems. It is thus necessary to create efficient quantum algorithms to solve high dimensional linear system of equations. Even without linearization, solving nonlinear partial differential equations (PDEs) would still demand, at the least, an efficient way to iteratively operate general non-unitary and non-hermitian matrices. In both cases, one needs a quantum algorithmic framework to perform iterative matrix-vector products and matrix inversion operations efficiently. They are generally of the form
[…]
where $M$ and $\mathbf{u}$ represent, respectively, a given matrix operator (e.g., discretized differential operator) and a complex-valued vector—e.g., a velocity field. Indeed, these operations are ubiquitous in computational applications in a wide range of fields.
In this paper, we propose algorithms for solving time-dependent PDEs by reducing them to the above matrix operations and marching them forward in time. In particular, we demonstrate the process by solving a fluid flow problem, including its implementation on an existing quantum computer, and make the case for its suitability on near-term machines. In the ensuing discussion, we first review the related literature and, with that background, highlight the advancement made by the current work. The proposed procedure opens up a promising new avenue towards realizing quantum acceleration of a wide range of practical problems beyond fluid dynamics.
I.1 Related Work
Quantum computing methods for PDEs, fluid flow problems included, may be classified into three major categories:
  • Category A – Quantum Direct Numerical Simulations (QDNS) [harrow2009quantum, berry2014high, berry2017quantum, subacsi2019quantum, liu2021efficient, childs2021high, krovi2023improved, an2023linear, fang2023time, bharadwaj2023hybrid, an2023quantum, berry2024quantum, brearley2024quantum, leong2022variational, lubasch2020variational, ingelmann2024two, wright2024noisy, pool2024nonlinear]
  • Category B – Lattice Boltzmann and Liouville equation methods [todorova2020quantum, budinski2021quantum, bakker2023quantum, li2023potential, itani2024quantum, succi2024three, succi2024ensemble, lin2022koopman, schalkers2024efficient, kocherla2023fully, kocherla2024two, penuel2024feasibility]
  • Category C – Schrödinger equation methods [jin2022quantum, jin2023quantum, lu2024quantum, meng2023quantum, jin2024quantum]
The present work belongs to category A, in which one solves nonlinear continuum equations without modelling or mapping to analogous problems. Below, we compare the present work with the state-of-the-art in categories B and C, whose hallmark is that they map the original problem into alternative formulations.
Category A – In this category, we consider the Quantum Linear Systems Algorithms (QLSA)—on which the present work is based—and Variational Quantum Algorithms (VQA). VQA approach solves the governing PDE as an optimization problem [leong2022variational, lubasch2020variational, ingelmann2024two, kocher2024numerical, pool2024nonlinear, bravo2023variational], where a parameterized velocity field is deemed to represent the desired solution when certain parameters correspond to the minimum of a suitably defined cost function. This approach has advantages such as relatively short depth circuits [cerezo2021variational, ingelmann2024two, kocher2024numerical], the possibility of encoding nonlinearity and minimal measurements (single-valued cost function output) [lubasch2020variational, pool2024nonlinear]. However, they are slow to converge [ingelmann2024two], and are prone to the barren-plateau problem, over-parameterization ansatzes [cerezo2021variational, holmes2022connecting, pool2024nonlinear], as well as the lack of complexity guarantees on quantum advantage [ingelmann2024two, leong2022variational].
QLSA offers complexity guarantees on quantum advantage and are devoid of barren-plateaus and heuristics from optimization. The QLSA-based methods entail solving a linear system of equations derived from the (non)linear governing PDEs. Such a construct is obtained by a combination of discretization methods (e.g., finite difference or finite element methods) and linearization techniques such as Carleman or Koopman embedding techniques, yielding a high dimensional linear system of equations, which is then solved using QLSA. Starting from the Harrow-Hassidim-Lloyd (HHL) algorithm [harrow2009quantum], these sets of algorithms have undergone a considerable evolution [berry2014high, berry2017quantum, subacsi2019quantum, liu2021efficient, childs2021high, krovi2023improved, an2023linear, bharadwaj2023hybrid, fang2023time, an2023quantum, berry2024quantum], by overcoming several challenges that impeded preceding proposals (see [aaronson2015read, bharadwaj2023hybrid]). This has been aided largely by advances made in improving Hamiltonian simulation algorithms [berry2007efficient, berry2014exponential, berry2015simulating, berry2015hamiltonian, low2017optimal, berry2020time, an2021time] as well as the introduction of the concept of Linear Combination of Unitaries [childs2012hamiltonian, childs2017quantum]. The central challenges of QLSA include (i) linear, non-optimal dependence on matrix system parameters such as grid size and number of time steps, matrix sparsity and condition number, and the specified accuracy; (ii) exponentially growing query complexity, requiring multiple copies of the initial state and repeated measurements and amplitude amplification of intermediate and final solution states; (iii) specific constraints on the properties of matrix operators such as hermiticity, unitarity, positive-definiteness and on the range of eigenvalues; (iv) stability requirements for time marching problems, and (v) the depth of quantum circuits due to quantum phase estimation methods and inefficient LCU decomposition. The more recent works, particularly those based on LCU, have addressed some of these aspects, while some [bharadwaj2023hybrid] have also (in the context of fluid dynamics) attempted to make the algorithms end-to-end, by addressing problems of state preparation and post-processing quantum information.
Category B – The Lattice Boltzmann Method offers an alternative, meso-scale formulation of fluid flows, where a discrete set of probability distributions are tracked on a fixed lattice, with a restricted set of velocity directions and magnitudes. Although the corresponding quantum algorithms exhibit certain innate advantages, they still face certain challenges [todorova2020quantum, budinski2021quantum, lin2022koopman, bakker2023quantum, li2023potential, kocherla2023fully, itani2024quantum, succi2024three, succi2024ensemble, schalkers2024efficient, kocherla2024two, penuel2024feasibility]. Besides the errors from inaccurate quantum circuit modelling of nonlinear collision-streaming operations [itani2024quantum] and physical boundary conditions [todorova2020quantum], their computational advantage tend to be downgraded because (a) qubit complexity scales linearly with problem size (grid size and time steps) when lattice positions are encoded as binary state vectors, leading to non-unitarity of streaming or collision steps [kocherla2023fully, schalkers2024importance], (b) repeated measurements are needed at every time step [kocherla2024two, budinski2021quantum] and (c) multiple copies are required of the initial state [itani2024quantum]. Although there have been efforts to ameliorate each of these constraints, a single algorithm that surpasses all these bottlenecks remains elusive. Another strategy would be to solve the linear Liouville equation [lin2022koopman, succi2024ensemble] that avoids the fundamentally nonlinear governing equation. However, this is still prone to errors from spurious effects that appear with increasing grid resolution. It seems that the best computational complexity for the algorithms in this category emerges when the problem is translated into a linear system of equations, solved subsequently by a Quantum Linear Systems Algorithm (QLSA) [li2023potential, penuel2024feasibility, bakker2023quantum] (category A).
Category C – Algorithms in this category map the original continuum governing PDEs into a Schrödinger or a Schrödinger-like equation. The algorithm presented in [meng2023quantum] maps the Navier-Stokes equations into an analogous hydrodynamic, nonlinear Schrödinger equation by using a Madelung transformation, which is then simulated as a quantum circuit. The upside of this approach stems from its innate quantum mechanical construction, but the overall complexity of the algorithm scales linearly with the problem size and an asymptotic computational advantage remains elusive. A potential alternative is that of Schrödingerisation [jin2022quantum, jin2023quantum, lu2024quantum, hu2024quantum, jin2024quantum], which offers a better computational complexity compared to other works within this category. The approach builds an exact mapping from a classical PDE into a dilated Schrödinger equation having an extra dimension with an effective Hamiltonian. However, in the absence of linearization methods, such mappings are restricted to a specific set of PDEs rather than to general nonlinear PDE’s such as the Navier-Stokes equations [lu2024quantum]. However, this approach tends to be a natural candidate for analog quantum computers than for current gate-based quantum computers, and exhibit a similar computational complexity as some earlier gate-based QLSA proposals. There is still a non-optimal, linear dependence (gate complexity) on the solution accuracy and time [hu2024quantum] for simulating problems that are similar to those considered in this paper.
I.2 Current Contribution
@@
This work presents a set of hybrid, quantum algorithms based on Linear Combination of Unitaries (LCU) for solving time-dependent PDEs, with an application to unsteady flows. We make use of a particular LCU decomposition, which has been used earlier in a different context of non-hermitian open quantum systems [schlimgen2021quantum] and quantum chemistry applications, and transform it into a broader framework of a hybrid quantum-classical algorithm for iterative matrix operations. This is accomplished by introducing six Time Marching Compact Quantum Circuit (TMCQC) algorithms to solve the time-dependent PDEs, admitting both explicit and implicit time stepping schemes. These algorithms require only two, or at most four, controlled unitaries for each LCU decomposition leading to relatively low-depth quantum circuits. We show that, of the six TMCQCs proposed, the gate complexity (of basic two- and one-qubit gates) is the best for (TMCQC-5,6), and is $\mathcal{O}\Big{(}s\epsilon,\log(N_{g}\tau),\textrm{polylog}(\epsilon/\varepsilon_{U},1/\varepsilon_{N}),\log^{-3}((\kappa-1)^{-1})\Big{)}$. Here, the governing PDEs are discretized using a finite difference method with $N_{g}$ as the grid size, $\tau$ as the number of time steps (thus an overall problem size of $\mathcal{O}(N_{g}\tau)$), $s$ and $\kappa$ being the sparsity and condition numbers of the matrix operators, $\epsilon$ a parameter controlling the LCU decomposition and $\varepsilon_{U}$ and $\varepsilon_{N}$ determine the accuracy of different approximations of the algorithm, as discussed in Section IV. We use these algorithms to simulate a linear advection-diffusion problem and use its analytical and classical solutions as reference for estimating the accuracy of the solutions obtained. The simplicity of this problem makes it an ideal candidate for assessing the performance of the quantum algorithm. We show that the results from the proposed algorithm agree qualitatively and quantitatively with analytical results.
The above complexity scaling is near-optimal in all parameters except $s$ and $\epsilon$. However, $s$ is equal to the order of finite difference scheme and therefore contributes only a small constant prefactor and $\epsilon$, as will shown in Section IV.6, may be maintained close to unity, thus adding only modest overhead to the complexity. There is thus the potential for an overall exponential advantage. The complexity (TMCQC-1,2,4) is either linear in $\tau$, or quadratic at worst, while still scaling optimally in other parameters. The qubit complexity scales as $\mathcal{O}(\log(N_{g}\tau))$ at best and $\mathcal{O}(\log(N_{g}\tau(\log(\varepsilon^{-1}_{N}))/\log((\kappa-1)^{-1})))$ at the worst, which is still optimal in both cases and provides an exponential memory advantage. These complexities are without quantum amplitude amplification, whose inclusion gives an improvement up to quadratic. These are significant advances with respect to existing proposals in categories B, C as well as VQA in category A.
Now with respect to the QLSA approaches, the improvements of the present algorithm are that it (a) avoids expensive Quantum Phase Estimation with a near-optimal dependence on the different parameters of the matrix operators; (b) admits non-unitary and non-hermitian matrices, without restrictions on the range of eigenvalues. In comparison with some recent, time-marching proposals [fang2023time, an2023quantum, berry2024quantum], it uses tools such as the Richardson extrapolation to lower the overall query complexity of the algorithm, thus ameliorating the dependence on the amplification ratio [fang2023time]; for the particular problem considered here, in comparison to [an2023quantum, fang2023time, brearley2024quantum], the present work has a complexity that is logarithmic in $\tau$, $\kappa$ (and $||M||$). (c) it provides complexity estimates for various parameters of the matrix, and avoids the need to compute integrals by using quadratures; and (e) it provides a simple and straightforward LCU decomposition that can be extended to other problems besides the specific example considered here.
The present end-to-end algorithm includes specific Quantum State Preparation and Quantum Post Processing algorithms to conserve quantum advantage, by introducing both explicit and implicit time stepping methods. We provide a fully gate level implementation of state vector simulations using QFlowS—an in-house, high-performance, quantum simulator [bharadwaj2023hybrid, bharadwaj2024qflows]. We also use a real quantum device (IBM Cairo) and study effects of noise using IBM Qiskit Aer platform and propose circuit designs and strategies for accurate, forward time simulations on near-term quantum devices. By choosing suitable problem sizes, our noisy simulation studies of the proposed quantum circuits indicate that the required gate error rates and execution times for convergent simulations are in the ballpark of the state-of-the-art devices and experimental specifications, thus hinting the possibility of retaining quantum advantage. The study also indicates how noise can be used to our advantage.
@@ II Governing equations
where the velocity varies only along y (wall-normal direction). We set $C=U$ (the constant advection velocity in the x-direction), and we put $D$, representing the diffusion coefficient, to unity. The initial condition $u(L/2,0)$ is chosen as the delta function $\delta(x)$ and the system is subject to a periodic boundary condition $u(0,t)=u(L,t)$. Such a setting also admits an analytical solution [ingelmann2024two], making it an ideal test-case to evaluate the accuracy of the quantum solutions. The algorithmic framework developed here is also agnostic to the linearity or otherwise of the PDEs. For instance, if we were to begin by considering the nonlinear case ($C=\mathbf{u}(x,t)$), a preceding linearization such as a Homotopy Analysis and Carlemann or Koopman method [liu2021efficient, giannakis2022embedding, lin2022koopman, bharadwaj2023quantum] can be applied to first obtain an approximate, higher dimensional linear system of equations. Such a system can then be solved using the proposed methods.
@@ IV.6 Query complexity and success probability
    does not diverge. If the time marching operator is norm-preserving (or the flow is statistically steady), then $||~{}|\Psi\rangle_{\tau}~{}||=||~{}|\Psi\rangle_{0}~{}||$ trivially. The alternative scenario of the norm-increasing flow is better, where $||~{}|\Psi\rangle_{\tau}~{}||\geq||~{}|\Psi\rangle_{0}~{}||$, corresponding to flows that experience constantly increasing net influx of energy. Both cases are in practice possible if we consider a constant forcing term in our governing PDEs. However, the more restrictive case is the viscous, dissipative flow whose norm decays with time; in this case the steady state of the flow corresponds to vanishingly small velocity fields due to constant dissipation of energy from the system. However, the advection-diffusion problem considered in this work is not as severe as can be seen from the analytical solution [ingelmann2024two]. The steady state corresponds to a constant, uniform velocity field with a small, yet non-zero norm (in this work, this ratio is = 1/2). For flows whose steady state norms are identically zero, the query complexity grows as $\mathcal{O}(||~{}|\Psi\rangle_{\tau}~{}||^{-1})$. In such cases, there could be two alternatives (which should be studied in more detail): (1) The problem of the growing pre-factor $\eta$ can be addressed by choosing a smaller $\epsilon$ such that $\eta^{G_{\textrm{L}}}\sim\mathcal{O}(||~{}|\Psi\rangle_{\tau}~{}||^{-1})$. (2) When this choice is not possible, one is limited to solving the problem for only short time horizons, for which the norm decays not too steeply.